OpenStudy (anonymous):

Show that 2-i and -2+i are square roots of 3-4i.

hartnn (hartnn):

let the square root of 3-4i be a+bi so, we have \(\large a+bi=\sqrt{3-4i}\) square both sides of this equation, what do u get ?

OpenStudy (anonymous):

\[a ^{2} + bi ^{2} = 3-4i ??\]

OpenStudy (anonymous):

\[(f+g)^2 ≠ f^2 +g^2\] \[(f+g)^2 = f^2 + 2fg +g^2\]